OpenStudy (anonymous):

When graphing y < 5x - 2, what kind of boundary line is used and where will the graph be shaded? 1. a broken line, shaded below the line or 2. a solid line, shaded below the line

OpenStudy (richyw):

if you have \(y<k\) then your values of \(y\) are everything less than but NOT equal to \(k\) and you would use a dotted line. If you have \(y\le k\) then \(y\) is everything less than or equal to \(k\) and you would use a solid line. If \(y<f(x)\), then you would shade all values under the function \(y=f(x)\). If \(y>f(x)\), then you would shade all values above the function \(y=f(x)\).